OpenStudy (anonymous):

Find the derivative of f(x) = - 9/x at x = -12.

OpenStudy (anonymous):

f(x) = -9 / x = -9x^-1 f'(x) = -9(-1)x^-2 = 9 / x^2 f'(-12) = 9 / (-12)^2 = 9 / 144 = 1 / 16

OpenStudy (anonymous):

My answer options are: 72/5, 5/72, 6/5,5/6..? @TURITW

OpenStudy (anonymous):

@Ahmad1 @oldrin.bataku ??

OpenStudy (anonymous):

Is the question correct?

OpenStudy (anonymous):

yeah, I tried doing it the way you did too, and I've got nothing..@TURITW

OpenStudy (anonymous):

@TURITW

OpenStudy (anonymous):

\[-9x ^{-1} = 9x^{-2} \rightarrow 9*144^{-1} =\frac{ 9 }{ 144 }=\frac{ 1 }{ 16 }\]

OpenStudy (anonymous):

^^^ doesn't work

OpenStudy (anonymous):

Probably question or answer is wrong

OpenStudy (anonymous):

well my teach wants me to find it by using the quotient property

OpenStudy (anonymous):

f(x) = -9 / x f'(x) = (x(0) - (-9)(1)) / x^2 = 9 / x^2

OpenStudy (anonymous):

nothing changes dude the answer still the same

OpenStudy (anonymous):

lol yeah so..i don't really know how to solve it because that's the only way i've used

OpenStudy (anonymous):

I got your point, you still using diffrence qutiont to find the derivative , just few weeks and you will be learning these rules